OpenStudy (anonymous):

Cos (-13pi/3) trigonomic functions

OpenStudy (jdoe0001):

hmmm what are you supposed to do? \(\bf cos(-\theta) = cos(\theta)\qquad thus\qquad cos\left(-\cfrac{13\pi}{3}\right) = cos\left(\cfrac{13\pi}{3}\right)\)